Andrea can ride her bike 2 mph faster than Ruby. It takes andrea 48 minutes less to travel 50 miles than it does Ruby. What is Andrea's rate in miles per hour?
Answer by mananth(16946)

Andrea can ride her bike 2 mph faster than Ruby.
Let Ruby's rate be x
Andrea's rate = x+2
distance = 50 miles
Ruby's time - Andrea's time = 48 minutes => 4/5 hours
50/x - 50/(x+2) = 4/5

50(x+2)-50x = x(x+2)*4/5
50x +100 -50x=(x^2+2x)4/5
100 = (4x^2 +8x)/5
500 = 4x^2 +8x
4x^2+8x-500=0
/4
x^2+2x-125 =0
x^2+2x+1 = 125+1
(x+1)^2 = 126
x+1 = +/- sqrt(126)
x= 10.22 mph
Andrea's speed = 12.22 mph